# Perils of Label Indeterminacy: A Case Study on Prediction of Neurological Recovery After Cardiac Arrest

### v1

@@ 4 Empirical Study
To assess the effects of label indeterminacy in practice, we conducted an empirical study in the healthcare domain, focused on predicting patient recovery after cardiac arrest. In this section, we begin by presenting the use case and a summary of the data used in our study (Section 4.1). In Section 4.2, we describe the complementary data of expert assessments collected via a panel [22]. Following this, we explore various modeling approaches to deal with unknown labels (Section 4.3) and compare these models’ performance against the portion of available ground-truth labels (Section 4.4). Finally, in Section 4.5, we compare the models’ predictions for cases with indeterminate labels.
4.1 Use Case and Historical Data
@@
Each year, over 150,000 Americans are hospitalized after resuscitation from cardiac arrest [45]. Many of these patients are initially in a coma and require life-sustaining therapies in the intensive care unit; around 70% do not survive their hospital stay [52, 9]. For comatose patients, the first major predictive task for doctors is typically to differentiate between patients who can awaken from coma and those for whom severe brain injury precludes awakening. In the US, the surrogate decision-makers for most patients who are judged by doctors to have a good chance of recovery opt to continue life-sustaining therapies to facilitate this recovery. Conversely, when doctors believe there is no chance of future awakening from coma, most decision-makers choose to Withdraw Life-Sustaining Therapies (WLST) in favor of a comfortable death. Awakening from coma after cardiac arrest can take several days, and identifying potential for awakening is challenging, especially early on  [61, 57, 52]. It is known that doctors are not perfectly accurate when making these prognostic judgments [67, 68]. Misguided therapeutic nihilism [30] and knowledge gaps contribute to WLST and avoidable deaths of patients who would have recovered had life-sustaining treatments been continued. Because death follows WLST decisions, recovery cannot be observed after WLST—this is illustrated in Figure 1.
@@
Figure 1: Obtaining labels depends on whether patients are withdrawn from life-sustaining therapies (WLST) or not. In the case of WLST, patients die, and we do not know whether they could have recovered had they not been withdrawn from life-sustaining therapies. For WLST cases, labels are unknown.
Given the complexity of neurological prognosis, there is substantial interest in leveraging AI-based decision support for predicting recovery potential [55, 31, 38, 47]. AI systems, including supervised machine learning, offer a prospect of improving prognostication after cardiac arrest by providing more accurate and timely outcome predictions [38]. The goal of most prognostic AI models is to predict recovery for individual patients based on their clinical characteristics. These models are trained using existing data and seek to guide future medical care. However, the presence of the selective labels problem poses the risk of creating harmful self-fulfilling prophecies, where flawed historical treatment decisions are repeated [14]. Hence, efforts have been made towards mitigating bias stemming from selective labels through statistical approaches [21].
@@
Data for our study are derived from a cohort of patients resuscitated from cardiac arrest who were hospitalized at a single US-based academic medical center. Clinical information includes patient demographics, details of the circumstances of cardiac arrest, daily clinical summaries (e.g., neurological exam findings, prognostic test results, laboratory tests) and recovery (yes/no). Recovery is defined as following verbal commands and surviving to hospital discharge. For our study, we consider a sample of 2,676 patients, in 1,026 of whom (38.3%) we observe the true outcome (either recovered or not) because therapies were not withdrawn, and in 1,650 of whom (61.7%) WLST results in unknowable labels. For the subset of WLST cases, we collected expert assessments to gauge their counterfactual recovery potential for the hypothetical case that they had been provided with ongoing therapies.
4.2 Panel Data
@@
An international panel of 38 leading experts in neurological prognostication was recruited via their professional reputation [22]. To be included in the panel, experts had to meet at least two of the following three criteria. They must have $(i)$ cared for over 150 post-arrest patients, $(ii)$ published more than three peer-reviewed articles on post-arrest care and prognostication in the past three years, and $(iii)$ played a leadership role in a major study of post-arrest care, such as a clinical trial.
@@
Over two years, IRB-approved one-hour case review sessions were conducted [22] with groups of three or more experts via Zoom. During each session, detailed clinical information was reviewed from individual patients who died after WLST. Every session included approximately four cases, depending on the complexity of the cases. Experts were presented with a structured clinical summary presentation and had an opportunity to ask clarifying questions. The experts then navigated to the data collection instrument and independently entered their counterfactual outcome estimate quantifying the likelihood of recovery if life-sustaining therapies had been continued. Experts estimated the counterfactuals along seven clinically relevant strata, depicted in Table 1.
Table 1: 7-point scale of expert assessments to predict the likelihood of patient recovery.
@@
4.3 Different Ways of Handling Unknown Labels
We outline the construction of ten predictive models, based on approaches to deal with indeterminate labels that have been used in the literature or that would be plausible to use. For each, we include a brief discussion of motivations and unverifiable assumptions. First, we establish some notation. Let $X$ be the set of clinical information used as covariates. Further, let $N$ be the set of patients with known labels $Y_{i}\in\{0,1\}$, $i\in N$, and let $W$ be the label indeterminacy set. For any WLST patient $j\in W$, we have three expert assessments, $e_{j}^{1},e_{j}^{2},e_{j}^{3}\in[0,1]$, [^1] We translate discrete expert ratings to soft labels that correspond to the upper bound of the respective rating category. For instance, a rating of 3 corresponds to a predicted likelihood of recovery of 0.1. which we treat as soft labels. [^2] Note that for two WLST patients $j$ and $k$, with $j\neq k$, the expert ratings $e_{j}^{\ell}$ and $e_{k}^{\ell}$, $\ell\in\{1,2,3\}$, do not generally come from the same expert. The notation and data setup is summarized in Figure 2.
@@ 4.3 Different Ways of Handling Unknown Labels
Figure 2: Data setup and taxonomy of our empirical study. The whole patient population is divided into patients with (blue, left) and without (orange, right) known labels. The subset of patients with known labels is $N$, and the label indeterminacy set (i.e., WLST patients) is $W$. For each WLST patient, we have three expert assessments, $e_{j}^{1},e_{j}^{2},e_{j}^{3}\in[0,1]$. Dashed borders indicate that the corresponding values are not ground truth.
4.3.1 Only Based on Historical Data
@@
A common approach to tackle the sampling bias in $M_{obs}$ is to apply inverse propensity weighting (IPW) during training [2, 41]. This is done by training a separate classifier that predicts the likelihood of observing the true label for each instance, and then weighing each instance $i\in N$ by their inverse probability of being observed. [^3] In our empirical experiments, we use logistic regression for estimating the propensity scores, with a test AUC of 0.8199. This way, instances that resemble WLST cases receive high weight during training, correcting the sampling bias. Note that in cases where the source of selection bias is historical expert decisions, this model is predicting those human decisions. Crucially, this approach assumes that all instances have a non-negligible probability of being part of the training sample—known as the positivity assumption—and it also assumes that the decision that determines whether a label is observed does not depend on any unobserved confounders. Both of these assumptions are hard to meet in practice [15]. The assumption of no confounders is not one that can be verified solely from the data, while the positivity assumption is testable but frequently found to be violated. In our data, neither of these assumptions holds; there are cases that have a negligible or null probability of non-withdrawal given that they present clear markers of poor neurological prognosis. Moreover, WLST decisions are affected by factors not contained in the data, such as information communicated by the family members, as well as medical information that is not part of the available covariates $X$.
@@
Instead of ignoring WLST cases, this approach, which is ubiquitous in biomedical research [14], imputes $\widetilde{Y}_{j}=0$ for all $j\in W$. It then uses all cases that underwent WLST in addition to the ones with known labels for training. The motivation for this approach is that WLST decisions are made by highly knowledgeable experts. Hence, these decisions are often thought to be a reliable label. While this approach ensures that WLST cases are represented, it risks erroneously associating factors that predict WLST with poor neurological outcome, because it assumes that historical decisions to withdraw life-sustaining therapies were correct. In particular, if the motivation to use machine learning is to improve the quality of decisions, it is contradictory to assume that historical decisions were perfect.
@@
An alternative way of estimating WLST labels is to impute them with labels from similar patients who were not withdrawn from life-sustaining therapies. A common method for achieving this is nearest-neighbor imputation [10]. In our study, we operationalize this for any $j\in W$ as follows: $\widetilde{Y}_{j}=Y_{i}$, with $i=\operatorname*{arg\,max}_{k}\left(\frac{X_{j}\cdot X_{k}}{\|X_{j}\cdot X_{k}\|}\right)$ for $k\in N$. This approach assumes knowledge of a relevant similarity metric, and for each WLST patient $j$, the presence of at least one patient $k\in N$ that is “similar enough.” The latter is equivalent to the positivity assumption, which we have discussed above. The former is also typically violated, as standard metrics, such as Euclidean distance, are unlikely to capture relevant similarity, and attempts to learn a task-specific metric encounter the same data issues as attempts to predict neurological recovery.
4.3.2 Leveraging Expert Assessments
@@
A prevalent practice in the literature on label aggregation [73] is to take the mean of multiple assessments to arrive at one label per patient. In our study, this means $\widetilde{Y}_{j}=\frac{1}{3}\cdot\left(e_{j}^{1}+e_{j}^{2}+e_{j}^{3}\right)$, $j\in W$. In many cases, this approach yields a model that is equivalent to $M_{exp_{all}}$. In this framing, however, useful information is removed from the data, and with it the potential to leverage that information by estimating quantities other than the most likely outcome. In particular, expert disagreement might be based on fundamentally different world views, in which case the “average expert” is a meaningless measure [44]. Moreover, by aggregating, any potentially insightful variance in expert responses gets lost.
@@
Instead of augmenting all WLST cases, the idea of this labeling approach is to only consider cases where experts agree in their assessments, and disregard those with substantial disagreement. The intuition behind this is that it may be desirable to learn from experts when they exhibit consistency, as this may be indicative of underlying expert consensus, while excluding assessments with observed disagreement—especially if in such cases we can default to observed outcomes [16]. We operationalize this in our study as follows:
@@
Augmenting only WLST cases where experts agree on the absence of recovery potential aims at mitigating the selective labels problem, which only affects our ability to observe the true outcome for WLST cases [15]. Thus, the intuition is that we may only want to rely on expert assessments when we do not have the possibility of observing a true outcome. We operationalize this as follows:
@@
This model combines the motivation of the previous model, $M_{exp_{agr_{w}}}$, with inverse propensity weighting (IPW) to address sampling bias [2, 41]. Here, it is important to note that cases where experts agree on WLST are precisely the ones that lead to a violation of the positivity assumption. Recall that the positivity assumption is violated when we never observe an outcome for certain types of patients, because based on current medical knowledge they are always withdrawn from life-sustaining therapies. When the panel agreement reflects this, incorporating these labels allows us to address the violation of this assumption. The approach first constructs the same WLST labels as $M_{exp_{agr_{w}}}$, treats them as ground truth, and then applies IPW during training. It relies on the same implicit, unverifiable assumptions about expert consistency as before.
4.4 Results: Predictions on Patients With Known Labels
@@
To evaluate the different approaches to deal with label indeterminacy introduced in Section 4.3, we first evaluate the performance of the respective models on patients for whom we have known labels, i.e., patients that did not undergo WLST. To do this, we set up a 5-fold cross-validation pipeline to make predictions on five disjoint holdout sets of patients with known labels. We use random forest regression models [6], which are commonly used in recovery prediction [47]. All ten models are trained on $(i)$ the cases with known labels (except for the holdout sets) plus $(ii)$ the respective WLST cases with estimated labels, if applicable (see Table 4 in Appendix A for a summary). As performance metric, we assess the area under the receiver operating characteristic (ROC) curve, short AUC, which is standard practice in medical diagnosis [34, 21]. The AUC score is a single-value measure between 0 and 1 that captures the predictive ability of a model. We also qualitatively assess the ROC curves themselves. ROC curves plot sensitivity (i.e., a model’s ability to correctly identify patients with no recovery potential) against 1–specificity (i.e., the share of patients that are incorrectly predicted to have no recovery potential) for different decision thresholds [33].
@@
First, we note that predictive performance in terms of AUC is similar to prior studies on recovery prediction with AI, which lends clinical validity to our results [21]. Importantly, in Table 2, we also see that the AUC scores are not significantly different between all ten models. Moreover, Figure 3 shows that sensitivity at different specificity levels is comparable across all models. In other words, the predictive performances of all ten models are nearly indistinguishable on the subset of patients with known labels. This shows that the choice of how to estimate and incorporate labels for WLST patients is arbitrary if we assess the different choices with respect to the resulting performance on patients with known labels. We have ten different ways of handling indeterminate labels, each with unverifiable or knowingly violated assumptions, and they all exhibit statistically comparable performance when evaluated with respect to the portion of cases for which we know the true outcome.
Table 2: Means and standard deviations (STD) of AUC scores from 5-fold cross validation on patients with known labels. AUC scores are values between 0 and 1; higher AUC scores indicate better predictive performance.
Mean 0.9247 0.9246 0.9249 0.9241 0.9269 0.9272 0.9260 0.9259 0.9244 0.9209
STD 0.0220 0.0243 0.0175 0.0236 0.0152 0.0172 0.0159 0.0144 0.0222 0.0217
@@
Figure 3: ROC curves with confidence bounds from 5-fold cross validation on patients with known labels. Shapes are indistinguishable across models.
4.5 Results: Predictions on Label Indeterminacy Set
A similar evaluation of predictions for patients with indeterminate labels is not possible because of the absence of ground truth. But we are able to analyze differences in predictions across the different models. On a holdout set of WLST cases, we can observe that the distributions of predictions for these patients vary significantly between models (see Figure 4). Intuitively, it seems plausible that the models which are only trained on known labels ($M_{obs}$ and $M_{obs+ip}$) issue relatively more optimistic predictions, because the reason we observe true labels for them is that experts in the past were optimistic enough to not withdraw life-sustaining therapies. Similarly, it appears sensible that the $M_{w_{corr}}$ model issues relatively more pessimistic predictions because it learns that all WLST patients have no chance of recovery—and conversely, cases in which we augment the data by explicitly selecting the most optimistic expert assessments may yield models that are more optimistic. Thus, based on the results shown in Figure 4, it would be possible that models only vary in their calibration while maintaining the same ranked ordering of patients. While this is typically not a big problem in machine learning, since probabilities can be calibrated via post-processing, in light of label indeterminacy, we cannot calibrate predictions because we have no grounds of knowing what the true outcome for WLST patients is. In other words, we cannot know whether optimistic or pessimistic predictions are justified.
@@ 4.5 Results: Predictions on Label Indeterminacy Set
Figure 4: Distribution of predictions on holdout sets of WLST cases. Distributions are strikingly different across models, but we have no way of knowing which predictions are “best” in light of label indeterminacy.
Importantly, in our study it is not the case that predictions are merely “shifted” up or down from one model to another; the ordering of patients also changes. In other words, the difference between the different models is not only explained by different calibration. For instance, there are 35 patients for whom the model $M_{exp_{max}}$ predicts a recovery potential of $>$10% and $M_{obs}$ only $\leq$1%. These pairwise comparisons are shown in Table 3 for all models, as well as in Tables 5 and 6 in Appendix B for more extreme discrepancies. Overall, 323 WLST patients (approx. 19.6% of our sample) have a predicted likelihood to recover of $\leq$1by at least one model and $>$10% by at least one other. [^4] The threshold of 1% is often used in clinical practice to inform decisions as to whether life-sustaining therapies should be continued or withdrawn [22]. When comparing ranks, [^5] We look at the ranking of patients to study whether the differences in predictions across models can be explained as a calibration issue, which we confirm is not the case. We acknowledge that the decision of whether to withdraw life-sustaining therapies is typically not a ranking problem—unless one is operating under limited resources, such as allocating respirators during the COVID-19 pandemic [74]. 74 patients are ranked in the bottom third by at least one model and in the top third by at least one other model in terms of their likelihood to recover. We visualize these stark differences in individual predictions in Figure 5 for two exemplary patients, where at least one model predicts a near-zero likelihood to recover and at least one other predicts $\gg$50%. In particular, the models providing optimistic versus pessimistic predictions can vary significantly depending on the patient. For instance, $M_{exp_{max}}$ estimates a recovery chance of over 70% for patient PUH-2015-150 (Figure 5 left), while $M_{nn}$ suggests a recovery likelihood of less than 5% for the same patient. In contrast, when we examine patient PUH-2019-009 (Figure 5 right), the predictions are reversed. Crucially, these differences—which have critical and potentially irreversible implications for patients—are entirely a result of unverifiable design choices made in handling indeterminate labels.

### v2

@@ 4. Empirical Study
To assess the effects of label indeterminacy in practice, we conducted an empirical study in the healthcare domain, focused on predicting patient recovery after cardiac arrest. In this section, we begin by presenting the use case and a summary of the data used in our study (Section 4.1). In Section 4.2, we describe the complementary data of expert assessments collected via a panel by Elmer et al. (2025). Following this, we explore various modeling approaches to deal with unknown labels (Section 4.3) and compare these models’ performance against the portion of available ground-truth labels (Section 4.4). Finally, in Section 4.5, we compare the models’ predictions for cases with indeterminate labels.
4.1. Use Case and Historical Data
@@
Each year, over 150,000 Americans are hospitalized after resuscitation from cardiac arrest (Martin et al., 2024). Many of these patients are initially in a coma and require life-sustaining therapies in the intensive care unit; around 70% do not survive their hospital stay (Nolan et al., 2021; Callaway et al., 2015). For comatose patients, the first major predictive task for doctors is typically to differentiate between patients who can awaken from coma (a necessary first step in recovery) and those for whom severe brain injury precludes awakening. In the US, the surrogate decision-makers for most patients who are judged by doctors to have a good chance of recovery opt to continue life-sustaining therapies to facilitate this recovery. Conversely, when doctors believe there is no chance of future awakening from coma, most decision-makers choose to Withdraw Life-Sustaining Therapies (WLST) in favor of a comfortable death. Awakening from coma after cardiac arrest can take several days, and identifying potential for awakening is challenging, especially early on  (Sandroni et al., 2014; Paul et al., 2016; Nolan et al., 2021). It is known that doctors are not perfectly accurate when making these prognostic judgments (Steinberg et al., 2020, 2022). Misguided therapeutic nihilism (Hemphill III and White, 2009) and knowledge gaps contribute to WLST and avoidable deaths of patients who would have recovered had life-sustaining treatments been continued. Because death follows WLST decisions, recovery cannot be observed after WLST—this is illustrated in Figure 1.
@@
Given the complexity of neurological prognosis, there is substantial interest in leveraging AI-based decision support for predicting recovery potential (Park et al., 2019; Hessulf et al., 2023; Kwon et al., 2019; Mayampurath et al., 2022). AI systems, including supervised machine learning, offer a prospect of improving prognostication after cardiac arrest by providing more accurate and timely outcome predictions (Kwon et al., 2019). The goal of most prognostic AI models is to predict recovery for individual patients based on their clinical characteristics. These models are trained using existing data and seek to guide future medical care. However, the presence of the selective labels problem poses the risk of creating harmful self-fulfilling prophecies, where flawed historical treatment decisions are repeated (De-Arteaga and Elmer, 2023). Hence, efforts have been made towards mitigating bias stemming from selective labels through statistical approaches (Elmer et al., 2023).
@@
Data for our study are derived from a cohort of patients resuscitated from cardiac arrest who were hospitalized at a single US-based academic medical center. Clinical information includes patient demographics, details of the circumstances of cardiac arrest, daily clinical summaries (e.g., neurological exam findings, prognostic test results, laboratory tests), and recovery (yes/no), which entails consciousness recovery (i.e., an ability to follow verbal commands) and survival to hospital discharge. For our study, we consider a sample of 2,676 patients, in 1,026 of whom (38.3%) we observe the true outcome (either recovered or not) because therapies were not withdrawn, and in 1,650 of whom (61.7%) WLST results in unknowable labels. For the subset of WLST cases, we collected expert assessments to gauge their counterfactual recovery potential for the hypothetical case that they had been provided with ongoing therapies.
4.2. Panel Data
@@
An international panel of 38 leading experts in neurological prognostication was recruited via their professional reputation (Elmer et al., 2025). To be included in the panel, experts had to meet at least two of the following three criteria. They must have $(i)$ cared for over 150 post-arrest patients, $(ii)$ published more than three peer-reviewed articles on post-arrest care and prognostication in the past three years, and $(iii)$ played a leadership role in a major study of post-arrest care, such as a clinical trial.
@@
Over two years, IRB-approved one-hour case review sessions were conducted (Elmer et al., 2025) with groups of three or more experts via Zoom. During each session, detailed clinical information was reviewed from individual patients who died after WLST. Every session included approximately four cases, depending on the complexity of the cases. Experts were presented with a structured clinical summary presentation and had an opportunity to ask clarifying questions. The experts then navigated to the data collection instrument and independently entered their counterfactual outcome estimate quantifying the likelihood of recovery if life-sustaining therapies had been continued. Experts estimated the counterfactuals along seven clinically relevant strata, depicted in Table 1.
Table 1. 7-point scale of expert assessments to predict the likelihood of patient recovery.
@@
4.3. Different Ways of Handling Unknown Labels
Figure 2. Data setup and taxonomy of our empirical study. The whole patient population is divided into patients with (blue, left) and without (orange, right) known labels. The subset of patients with known labels is $N$, and the label indeterminacy set (i.e., WLST patients) is $W$. For each WLST patient $j\in W$, we have three expert assessments, $e_{j}^{1},e_{j}^{2},e_{j}^{3}\in[0,1]$. Dashed borders indicate that the corresponding values are not ground truth.
[…]
Data setup and taxonomy of the case study. For some patients, there are ground-truth labels; for others (WLST), there are three expert assessments.
We outline the construction of ten predictive models, based on approaches to deal with indeterminate labels that have been used in the literature or that would be plausible to use. For each, we include a brief discussion of motivations and unverifiable assumptions. First, we establish some notation. Let $X$ be the set of clinical information used as covariates. Further, let $N$ be the set of patients with known labels $Y_{i}\in\{0,1\}$, $i\in N$, and let $W$ be the label indeterminacy set. For any WLST patient $j\in W$, we have three expert assessments, $e_{j}^{1},e_{j}^{2},e_{j}^{3}\in[0,1]$, [^1] We translate discrete expert ratings to soft labels that correspond to the upper bound of the respective rating category. For instance, a rating of 3 corresponds to a predicted likelihood of recovery of 10%. which we treat as soft labels. [^2] Note that for two WLST patients $j$ and $k$, with $j\neq k$, the expert ratings $e_{j}^{\ell}$ and $e_{k}^{\ell}$, $\ell\in\{1,2,3\}$, do not generally come from the same expert. The notation and data setup is summarized in Figure 2.
@@ 4.3. Different Ways of Handling Unknown Labels
4.3.1. Only Based on Historical Data
@@
A common approach to tackle the sampling bias in $M_{obs}$ is to apply inverse propensity weighting (IPW) during training (Austin and Stuart, 2015; Little and Rubin, 2019). This is done by training a separate classifier that predicts the likelihood of observing the true label for each instance, and then weighing each instance $i\in N$ by their inverse probability of being observed. [^3] In our empirical experiments, we use logistic regression for estimating the propensity scores, with a test AUC of .8199. This way, instances that resemble WLST cases receive high weight during training, correcting the sampling bias. Note that in cases where the source of selection bias are historical expert decisions, this model is predicting those human decisions. Crucially, this approach assumes that all instances have a non-negligible probability of being part of the training sample—known as the positivity assumption—and it also assumes that the decision that determines whether a label is observed does not depend on any unobserved confounders. Both of these assumptions are hard to meet in practice (De-Arteaga et al., 2018). The assumption of no confounders is not one that can be verified solely from the data, while the positivity assumption is testable but frequently found to be violated. In our data, neither of these assumptions hold; there are cases that have a negligible or null probability of non-withdrawal given that they present clear markers of poor neurological prognosis. Moreover, WLST decisions are affected by factors not contained in the data, such as information communicated by the family members, as well as medical information that is not part of the available covariates $X$.
@@
Instead of ignoring WLST cases, this approach, which is ubiquitous in biomedical research (De-Arteaga and Elmer, 2023), imputes $\widetilde{Y}_{j}=0$ for all $j\in W$. It then uses all cases that underwent WLST in addition to the ones with known labels for training. The motivation for this approach is that WLST decisions are made by highly knowledgeable experts. Hence, these decisions are often thought to be a reliable label. While this approach ensures that WLST cases are represented, it risks erroneously associating factors that predict WLST with poor neurological outcome, because it assumes that historical decisions to withdraw life-sustaining therapies were correct. In particular, if the motivation to use machine learning is to improve the quality of decisions, it is contradictory to assume that historical decisions were perfect.
@@
An alternative way of estimating WLST labels is to impute them with labels from similar patients who were not withdrawn from life-sustaining therapies. A common method for achieving this is nearest-neighbor imputation (Chen and Shao, 2000). In our study, we operationalize this for any $j\in W$ as follows: $\widetilde{Y}_{j}=Y_{i}$, with $i=\operatorname*{arg\,max}_{k}\left(\frac{X_{j}\cdot X_{k}}{\|X_{j}\cdot X_{k}\|}\right)$ for $k\in N$. This approach assumes knowledge of a relevant similarity metric, and for each WLST patient $j$, the presence of at least one patient $k\in N$ that is “similar enough.” The latter is equivalent to the positivity assumption, which we have discussed above. The former is also typically violated, as standard metrics, such as Euclidean distance, are unlikely to capture relevant similarity, and attempts to learn a task-specific metric encounter the same data issues as attempts to predict neurological recovery.
4.3.2. Leveraging Expert Assessments
@@
A prevalent practice in the literature on label aggregation (Wei et al., 2023) is to take the mean of multiple assessments to arrive at one label per patient. In our study, this means $\widetilde{Y}_{j}=\frac{1}{3}\cdot\left(e_{j}^{1}+e_{j}^{2}+e_{j}^{3}\right)$, $j\in W$. In many cases, this approach yields a model that is equivalent to $M_{exp_{all}}$. In this framing, however, useful information is removed from the data, and with it the potential to leverage that information by estimating quantities other than the most likely outcome. In particular, expert disagreement might be based on fundamentally different worldviews, in which case the “average expert” is a meaningless measure (Manski et al., 2016). Moreover, by aggregating, any potentially insightful variance in expert responses gets lost.
@@
Instead of augmenting all WLST cases, the idea of this labeling approach is to only consider cases where experts agree in their assessments, and disregard those with substantial disagreement. The intuition behind this is that it may be desirable to learn from experts when they exhibit consistency, as this may be indicative of underlying expert consensus, while excluding assessments with observed disagreement—especially if in such cases we can default to observed outcomes (De-Arteaga et al., 2021). We operationalize this in our study as follows:
@@
Table 2. Means and standard deviations (STD) of AUC scores from 5-fold cross validation on patients with known labels. AUC scores are values between 0 and 1; higher AUC scores indicate better predictive performance.
Mean .9247 .9246 .9249 .9241 .9269 .9272 .9260 .9259 .9244 .9209
STD .0220 .0243 .0175 .0236 .0152 .0172 .0159 .0144 .0222 .0217
@@
Augmenting only WLST cases where experts agree on the absence of recovery potential aims at mitigating the selective labels problem, which only affects our ability to observe the true outcome for WLST cases (De-Arteaga et al., 2018). Thus, the intuition is that we may only want to rely on expert assessments when we do not have the possibility of observing a true outcome. We operationalize this as follows:
@@
This model combines the motivation of the previous model, $M_{exp_{agr_{w}}}$, with inverse propensity weighting (IPW) to address sampling bias (Austin and Stuart, 2015; Little and Rubin, 2019). Here, it is important to note that cases where experts agree on WLST are precisely the ones that lead to a violation of the positivity assumption. Recall that the positivity assumption is violated when we never observe an outcome for certain types of patients because, based on current medical knowledge, they are always withdrawn from life-sustaining therapies. When the panel agreement reflects this, incorporating these labels allows us to address the violation of this assumption. The approach first constructs the same WLST labels as $M_{exp_{agr_{w}}}$, treats them as ground truth, and then applies IPW during training. It relies on the same implicit, unverifiable assumptions about expert consistency as before.
4.4. Results: Predictions on Patients With Known Labels
@@
To evaluate the different approaches to deal with label indeterminacy introduced in Section 4.3, we first evaluate the performance of the respective models on patients for whom we have known labels, i.e., patients that did not undergo WLST. To do this, we set up a 5-fold cross-validation pipeline to make predictions on five disjoint holdout sets. We use random forest regression models (Breiman, 2001), which are commonly used in recovery prediction (Mayampurath et al., 2022). All ten models are trained on $(i)$ the cases with known labels (except for the holdout sets) plus $(ii)$ the respective WLST cases with estimated labels, if applicable (see Table 4 in Appendix A for a summary). As performance metric, we assess the area under the receiver operating characteristic (ROC) curve, short AUC, which is standard practice in medical diagnosis (Huang and Ling, 2005; Elmer et al., 2023). The AUC score is a measure between 0 and 1 that captures the predictive ability of a model. We also qualitatively assess the ROC curves themselves. ROC curves plot sensitivity (i.e., a model’s ability to correctly identify patients with no recovery potential) against 1–specificity (i.e., the share of patients that are incorrectly predicted to have no recovery potential) for different decision thresholds (Hoo et al., 2017).
@@
First, we note that predictive performance in terms of AUC is similar to prior studies on recovery prediction with AI, which lends clinical validity to our results (Elmer et al., 2023). Importantly, in Table 2, we also see that the AUC scores are not significantly different between all ten models. Moreover, Figure 3 shows that sensitivity at different specificity levels is comparable across all models. In other words, the predictive performances of all ten models are nearly indistinguishable on the subset of patients with known labels. This shows that the choice of how to estimate and incorporate labels for WLST patients is arbitrary if we assess the different choices with respect to the resulting performance on patients with known labels. We have ten different ways of handling indeterminate labels, each with unverifiable or knowingly violated assumptions, and they all exhibit statistically comparable performance when evaluated with respect to the portion of cases for which we know the true outcome.
@@
Figure 3. ROC curves with confidence bounds from 5-fold cross validation on patients with known labels. Shapes are nearly indistinguishable across models.
[…]
AUC curves for each model evaluated on patients with known labels. Shapes are nearly indistinguishable across models.
@@
Figure 4. Distribution of predictions on holdout sets of WLST cases. Distributions are strikingly different across models, but we have no way of knowing which predictions are “best” in light of label indeterminacy.
[…]
Distribution of predictions for WLST patients by model. Distributions are strikingly different.
4.5. Results: Predictions on Label Indeterminacy Set
A similar evaluation is not possible for patients with indeterminate labels because of the absence of ground truth. But we are able to analyze differences in predictions across the different models. Assuming a threshold of 1% recovery likelihood—which is often used in clinical practice to inform decisions as to whether life-sustaining therapies should be continued or withdrawn (Elmer et al., 2025)—the following percentages of WLST patients would remain withdrawn under each model: 30.1% ($M_{obs}$), 31.0% ($M_{obs+ip}$), 52.0% ($M_{w_{corr}}$), 23.4% ($M_{nn}$), 6.7% ($M_{exp_{all}}$), 6.9% ($M_{exp_{avg}}$), 0.6% ($M_{exp_{max}}$), 10.9% ($M_{exp_{agr}}$), 43.9% ($M_{exp_{agr_{w}}}$), 42.7% ($M_{exp_{agr_{w}}+ip}$). These stark differences indicate that label indeterminacy can have crucial practical implications. We also see in Figure 4 that the overall distributions of predictions for WLST patients vary significantly between models.
Intuitively, it seems plausible that the models which are only trained on known labels ($M_{obs}$ and $M_{obs+ip}$) may issue relatively more “optimistic” predictions, because the reason we observe true labels for them is that experts in the past were optimistic enough to not withdraw life-sustaining therapies. Similarly, it appears sensible that the $M_{w_{corr}}$ model may issue relatively more pessimistic predictions if it learns that all WLST patients have no chance of recovery—and conversely, cases in which we augment the data by explicitly selecting the most optimistic expert assessments may yield models that are more optimistic. Thus, based on the results shown in Figure 4, it would be possible that models only vary in their calibration while maintaining the same ranked ordering of patients. While this is typically not a big problem in machine learning, since probabilities can be calibrated via post-processing, in light of label indeterminacy, we would not be able to calibrate predictions because we have no grounds for knowing what the true outcome for WLST patients is.